OpenStudy (teamus1021):

The first four terms of a sequence are shown below. 8, 5, 2, -1 Which of the following functions best defines this sequence? f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (teamus1021):

thank u for helping me

jigglypuff314 (jigglypuff314):

@Teamus1021 You can try finding the values of f(2), f(3), and f(4) to try to see which equation matches what you got

OpenStudy (teamus1021):

i dont know much about algebra can u break it down more for me im just learning

OpenStudy (teamus1021):

would it be the 1st one

jigglypuff314 (jigglypuff314):

Of course, we can do this step by step ^_^ I think we should try the process of ellimination first, we are given that f(1) = 8 for every choice then let's take A for example f(n + 1) = f(n) + 3 if n = 1 then f(2) = f(1) + 3 do you see how I got that? :)

OpenStudy (teamus1021):

yes alittle bit

jigglypuff314 (jigglypuff314):

the idea is that we want to know if f(2) = 5 or not with the equation from A f(2) = f(1) +3 we plug in f(1) = 8 to get f(2) = 8 + 3 right? :)

OpenStudy (teamus1021):

yes

OpenStudy (teamus1021):

o its the last one we have to subtract to get 5 right

jigglypuff314 (jigglypuff314):

Perfect! You got it :D well done ^_^

OpenStudy (teamus1021):

thanz

OpenStudy (teamus1021):

i have one more can u help

jigglypuff314 (jigglypuff314):

sure thing ^_^ make a new question and tag me there :)

OpenStudy (teamus1021):

An airplane ascends at the rate of 40 feet per minute. The function below shows the height of the airplane from the ground f(t) after t minutes. f(t) = 40t + 520 Which statement best describes the y intercept of the graph of the function?

OpenStudy (teamus1021):

It is equal to 520 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 560 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 40 and represents the height of the airplane above the ground, in feet, before it starts ascending. It is equal to 520 and represents the height of the airplane above the ground, in feet, before it starts ascending.

jigglypuff314 (jigglypuff314):

it wants us to examine the "y intercept" of the given equation it is in the y intercept form y = mx + b where b is the y intercept

jigglypuff314 (jigglypuff314):

so we should be looking for the answer choice that involves the number 520 because 520 is the y intercept in the equation f(t) = 4t + 520